how to find the othrocenter of a trianle whoose vertices are A(0,0) b(-1,3) c (2,-1)

how  to  find   the  othrocenter  of  a  trianle whoose  vertices  are  A(0,0)  b(-1,3)  c (2,-1)

Grade:

1 Answers

Badiuddin askIITians.ismu Expert
13 years ago

Dear gaurav

Orthocenter is the point of intersection of altitude

so find the equation of line which is perpendicular to the line joing point (-1,3) and (2,-1) and passes through (0,0)

slope of the line BC  = -1/2

slope of the line perpendicular to it =2

so equation of line :   y-3 =2 (x+1)

                          or  2x-y +5=0

 

now find another line which is perpendicular to the line joing point (0,0) and (2,-1) and passes through (-1,-3)

slope of the line AC  = -4/3

slope of the line perpendicular to it =3/4

so equation of line :   y =3/4 x

                          or  3x-4y =0

 

solve these two equation

x=-4  , y =-3

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed solution very quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin